OpenStudy (anonymous):

the sum of the digits of a two- digit number is 9. the sum of the number and its units digits is 74. find the number

OpenStudy (asnaseer):

call the digits of this 2 digit number t and u where t is the tens digit and u is the units digit. then, from the information given to you, you have: t + u = 9 10t + u + u = 74 solve this system of equations to find t and u

OpenStudy (anonymous):

ok i see what you did

OpenStudy (asnaseer):

gr8!

OpenStudy (anonymous):

@asnaseer do i solve each equation?

OpenStudy (anonymous):

or do i solve the equations together

OpenStudy (asnaseer):

you need to solve the "system" of equations, i.e. solve them together. one way is to use the first equation to find an expression of t in terms of u, and substitute that into the second equation and solve for u. then find t from the first equation once you know u.

OpenStudy (anonymous):

ok can u help me to solve it step by step please

OpenStudy (asnaseer):

why don't you make a start and I'll correct you if you start going wrong anywhere?

OpenStudy (anonymous):

ok thanks

OpenStudy (asnaseer):

np :)

OpenStudy (anonymous):

can u get me started....i have no idea how to start :(

OpenStudy (asnaseer):

ok, we have the following equation to solve: t + u = 9 10t + u + u = 74 so first get an expression for t from the first equation - what do you get?

OpenStudy (anonymous):

do i add t+10t???

OpenStudy (asnaseer):

look at this equation: t + u = 9 rearrange it to get it into the form: t = ...

OpenStudy (anonymous):

9?

OpenStudy (asnaseer):

t + u = 9 we need to get u from the left-hand-side of this equation to the right-hand-side. we do this by subtracting u from both sides of the equation. what would we end up with after doing that?

OpenStudy (anonymous):

so do i do (u-)t+(u-)u=9

OpenStudy (asnaseer):

ok - it looks like you first need to learn some of the fundamentals of algebra. I would advise you to take a look through some of the short videos on this site: http://www.khanacademy.org/math/algebra/solving-linear-equations they are very good and will help you get up to speed with this.